### v1

@@ 4.1 Data Availability
We use three main data sources in this work: the CoVaxxy Tweets database, which contains Tweets related to COVID vaccines covaxxy and is available at doi.org/10.5281/zenodo.4526494; CDC records of COVID cases, deaths, and vaccinations in US counties CDC2021 ; and Mønsted and Lehmann’s antivaccine tweets dataset monstedAlgorithmicDetectionAnalysis2019 , which supplements our own labeled data for training a tweet classifier. Details are provided in Supplementary Methods, and data is available at github.com/osome-iu/effects_of_antivax_tweets_on_covid19_outcomes.
@@ 4.2 Antivaccine Tweet Classifier
To track the volume of antivaccine content produced in each county, we built a text classifier that determines if a tweet expresses antivaccine sentiment. The classifier takes the text of a tweet as input and returns a label, either “antivaccine” or “other.” It was trained on 4,200 tweets labeled by two human annotators and 2,000 tweets labeled by Mønsted and Lehmann monstedAlgorithmicDetectionAnalysis2019 . The classifier achieved good accuracy ($F_{1}=0.74$) on a hold-out set of 900 tweets. Further details about the data annotations, classification model, and results are found in Supplementary Methods.
@@ 4.3 Exposure to Antivaccine Tweets
To measure the impact of antivaccine tweets on people’s propensity to get vaccinated, we measure the amount of antivaccine Twitter content to which people are exposed at the county level. Intuitively, if a population is strongly connected to other populations that produce a lot of antivaccine content, then its exposure is high. The per-capita antivaccine exposure rate in county $i$ at time $t$ is defined as
@@
where $N_{i}$ is the population of county $i$, $T_{j,t}$ is the number of antivaccine tweets in county $j$ during time window $t$, and $W_{ij}$ is the number of times that COVID-related tweets posted by users in county $j$ were retweeted by users in county $i$ during our observation period.
@@ 4.4 SIRVA Model
In the SIRVA epidemic model (see Figure 2), we assume that vaccinated people ($V$) do not become infected and that vaccine-hesitant people ($A=\alpha S$) do not become vaccinated. The dynamic equations for each county can be written as:
@@
Finally, we break the conversion rate $\gamma$ into components due to exposure to antivaccine tweets ($\gamma_{e}$) and other factors ($\gamma_{p}$): $\gamma=\gamma_{p}+\gamma_{e}E_{t}$, where $E_{t}$ is the antivaccine exposure rate in the county at time $t$ (specifically, the exposure over the previous eight days).
We also define a simpler comparison model, SIRV, which is a special case of SIRVA where $\alpha_{t}=0$ for all $t$. This allows us to test how including vaccine hesitancy and its dynamics impacts the model’s predictive performance.
Another possible model would include a feedback term, $\gamma_{a}\alpha$, in $\gamma$, accounting for the spread of vaccine hesitancy within a county as a social contagion. We also tested a model with this dynamic effect and found similar results to the SIRVA model described here. We dropped this feature of the model for simplicity.
4.5 Effect Estimations
We use causal graphical modeling (see Supplementary Methods) to derive an expression for the average treatment effect (ATE) of exposure $E$ on the vaccine uptake rate $\dot{V}=\frac{dV}{dt}$ from the the dynamical equations of the SIRVA model (Eq. 27):
$\text{ATE}=\frac{d}{dE_{t-1}}\mathbb{E}[\dot{V_{t}}\mid\alpha_{t-1}]\approx-\mathbb{E}[\nu_{t-1}\,S_{t-1,i}\;\gamma_{e}\,(1-\alpha_{t-1,i})]$ (8)
where the expectation value is taken over times $t$, counties $i$, and the posteriors of $\nu$, $\gamma_{e}$, $\alpha$.
Given the ATE from Eq. 8, we wish to calculate how many vaccinations were prevented in each county and nationwide. Assuming the total change in vaccine uptake rate is relatively small, we can estimate the number of vaccinations prevented each day in each county as ${\Delta V}_{t,i}\approx(\text{ATE})E_{t-1,i}$, and then use this estimation to calculate the total number of vaccines prevented nationwide as ${\Delta V}=N\,\Delta t\;\mathbb{E}_{i,t}[\frac{\Delta V_{t,i}}{N_{i}}]$, where $N$ is the total population of the US and $\Delta t$ is the length of the observation period (see Supplementary Methods).
After calculating the number of people across the whole US who remained unvaccinated as a result of exposure to antivaccine tweets (750,000), we wish to estimate how many cases and deaths may have been prevented among these people if they had been vaccinated instead. Assuming their infection and death rates were typical for unvaccinated people during the February-August 2021 time period (3,870 cases and 57 deaths per 100,000 people, see Supplementary Methods), we estimate that about 29,000 cases and 430 deaths occurred in this population. Of these cases and deaths, approximately 93% of cases and 94% of deaths may have been prevented by vaccination cdc_covid_vaccine_effectiveness ; Bernal2021_covid_efficacy , yielding the numbers of COVID cases and deaths attributable to antivaccine tweets, as reported in our results.
4.6 Model Selection Criteria
We can measure each model’s expected out-of-sample model performance using leave-one-out cross validation (LOO), approximated with Pareto-smoothed importance sampling (PSIS). The PSIS-LOO criterion vehtariPracticalBayesianModel2017 is designed to estimate out-of-sample predictive performance by approximating the expected log pointwise predictive density (ELPD) for a new dataset from the observed dataset without refitting the model to the data. The criterion is robust and efficient and represents the current state of the art for Bayesian model comparison. We therefore compare the performance of the SIRVA model to simpler models by measuring each model’s Bayesian LOO estimate of ELPD (ELPD-LOO). Values of ELPD-LOO closer to zero indicates better fit to the data. Metrics were computed on a sampled data set of 400 random counties and 24 evenly spaced dates spanning the observation period from February to August 2021.
@@ 4.6 Model Selection Criteria
Supplementary information contains supplementary methods and supplementary figures 1–2.
[…]
Acknowledgments We are grateful to Marissa Donofrio for annotating tweets, to Bjarke Mønsted and Sune Lehmann for sharing additional annotation data, and to Vincent Jansen, Alessandro Flammini, and YY Ahn for helpful discussion. This work was supported in part by the Knight Foundation and Craig Newmark Philanthropies.
@@
The authors declare no competing interests.
@@ A.1 Data
The primary data source for this work is the CoVaxxy project covaxxy from Indiana University’s Observatory on Social Media. CoVaxxy collects tweets related to COVID vaccines and vaccine hesitancy, and geolocates the tweets to US counties when possible. Using this dataset, we can track the online discourse surrounding COVID vaccination in individual US counties. In this study we used data from February 6th to August 9th, 2021, and used only tweets geolocated to US counties. The beginning of this window coincides with vaccines becoming widely available in the US, and the end of the window marks approximately the time when the vaccine uptake rate began to slow substantially. The primary features of this data are the text content, timestamps, and geolocations of the tweets.
The US Center for Disease Control (CDC) collects and publishes data on COVID health outcomes and vaccination in US counties, including cumulative cases, deaths, and vaccinations CDC2021 for each county, for each day. This is the source of the public health metrics in this work. Due to differences in local reporting systems and reporting schedules, some entries in the cumulative counts are imputed by CDC as the last known value until they are updated with new reports from the counties or states. Counties within the state of Texas are excluded from our dataset because the official CDC data on COVID vaccinations does not contain information about Texas counties until after October 22, 2021.
To train our antivaccine tweet classifier, we require a dataset of tweets labeled as “antivaccine” or “other.” We supplement our own labeled data with a similar dataset created by Mønsted and Lehmann monstedAlgorithmicDetectionAnalysis2019 . Restrictions apply to the availability of this dataset, which therefore is not publicly available. The data is, however, available from the authors upon reasonable request. Although this dataset is not specific to COVID antivaccine sentiment, we found its content to be similar enough to improve the performance of our classifier. We used the Mønsted and Lehmann dataset for training our model but not as part of test data used to evaluate our classifier’s performance.
@@ A.2 Antivaccine Tweet Classifier
To track the volume of antivaccine content produced in each county, we built a text classifier that determines if a tweet is expressing antivaccine sentiment. The classifier takes the text of a tweet as input and returns a label, either “antivaccine” or “other.” The classifier is a neural network based on the RoBERTa language model liu2019roberta , and is trained on a set of manually labeled tweets.
Our labeled training and test data were coded by two human annotators and examined for agreement. Cases where annotators disagreed were discarded. To be labeled as antivaccine, a tweet must express the belief that safe, effective COVID vaccines are bad, ineffective, not actually a vaccine, or harmful (without specific evidence). Most tweets expressing the belief that COVID vaccines are harmful made one of a few common claims, so these common claims were manually checked using reputable fact-checkers (e.g., PolitiFact, FactCheck.org) and CDC publications; in general, the common claims of harm were found to be false and labeled as antivaccine.
The classifier was trained on 4,200 labeled tweets from the CoVaxxy dataset. To increase training data volume and variety, we also added to this training set 2,000 tweets randomly selected from the Mønsted and Lehmann dataset monstedAlgorithmicDetectionAnalysis2019 , which were labeled by three human annotators; cases where annotators disagreed were discarded. The model was evaluated on tested on a hold-out set of 900 labeled tweets from the CoVaxxy dataset, labeled by the same method and annotators as the training data. We did not use cross-validation because (i) the training data includes Mønsted and Lehmann’s tweets, which do not match 2021 COVID-related tweets, and (ii) our sample of training tweets from the CoVaxxy dataset deliberately included a greater proportion of antivaccine tweets to help with classifier learning. Instead we produced the test dataset by pure random sampling from the CoVaxxxy data to ensure that classification metrics like $F_{1}$ are unbiased.
The classifier was evaluated using common classification metrics. It has an accuracy of $0.94\pm 0.02$, an $F_{1}$ of $0.74\pm 0.07$, and a Matthews Correlation Coefficient chiccoMatthewsCorrelationCoefficient2021 of $0.72\pm 0.08$, where 95% confidence intervals are computed by the bootstrap procedure efronIntroductionBootstrap1993 . This classifier was used to determine the number of antivaccine tweets geolocated in each county on each date. While some individual tweets may be misclassified, we believe the performance of the classifier is adequate to determine population-level trends and relative magnitudes in the prevalence of antivaccine tweets.
@@ A.3 Estimating Parameters
Our goal is to infer the parameters of the SIRVA model from the data. In particular, we want to infer $\gamma_{e}$ to understand whether exposure to antivaccine tweets leads additional people to become unwilling to be vaccinated. To estimate the likely range of the model parameters, we use Bayesian inference with MCMC sampling hamra_markov_2013 . The goal of Bayesian inference is to find the probability distribution (which we call the posterior distribution) that describes the likely values of a parameter given the data. From the posterior distributions, we can get a mean estimate, the 95% high probability interval (“HDI”), and a $p$-value for each parameter.
At a high level, Bayesian MCMC inference samples random parameter values proportionally to their likelihood given the data. The MCMC algorithm begins by sampling from a prior distribution, which is defined by specifying plausible ranges of the parameters, and gradually converges to the posterior distribution. We use the standard NUTS algorithm hoffmanNoUTurnSamplerAdaptively implemented in the NumPyro package phan2019composable .
To specify a likelihood function for the SIRVA model given the data, we will define probability distributions for the daily changes in each of our observed variables: cumulative cases ($C=I+R$), vaccinations ($V$), susceptible individuals ($S=N-(V+C)$), and recovered individuals ($R$). We assume $R_{t}\approx C_{t-8}$, based on a typical time from initial symptoms to non-infectious status of 10 days cdc_isolation_duration , and a lag time between initial symptoms and a positive test of about 2 days.
@@
These distributions have high variance relative to the plausible parameter ranges so that the priors do not strongly influence the final inferred parameter values. For the vaccine-hesitancy parameters, $\gamma_{e}$ and $\gamma_{p}$, we choose priors centered at zero. This way, if the posterior is found to be non-zero, we know the prior did not bias that conclusion. For $\alpha_{0}$, we choose a weak prior based on general estimates of vaccine hesitancy in the population aspe_vaccine_hesitancy . We use these priors:
@@
In our model, the parameters $\nu$ and $\beta$ can vary over time, following prior work dehning_inferring_2020 ; marinov_adaptive_2022 . The $\nu$ parameter is allowed to change once every six weeks to account for changing nationwide vaccine availability. The $\beta$ parameter is allowed to change once every three weeks to account for changing mean reproduction numbers associated with the emergence of new variants (e.g., the Delta variant late in our observation window) and changing public health policies (e.g., the imposition or lifting of lockdowns and mask mandates). The multiple values of these parameters are inferred by the MCMC sampler just like the other parameters.
Our dataset comprises 1,319 counties and 188 dates for which we have sufficient Twitter and public health data. We use a subsample of this data to perform our inferences. We sample by date to minimize the effect of temporal autocorrelations in the data; specifically, we use every 8th day in the time series data for each county. This interval is chosen for two reasons: (i) it’s slightly more than a week, so it smooths over data lags associated with weekly reporting cycles, and (ii) it is the approximate recovery time used to compute $R$. We also randomly sample by county to limit the number of model parameters (each county $i$ introduces an additional parameter $\alpha_{0,i}$). Our Bayesian MCMC inference tools struggle with numerical stability when the numbers of parameters and data points get too large. We therefore use 400 randomly selected counties and 24 evenly-spaced dates from our data. Although this sampling ultimately reduces the precision of our inferred parameter values, we do not believe it introduces any systematic biases; results are consistent across different random samples of the counties.

### v2

@@ 5.1 Data availability
We use three main data sources in this work. The CoVaxxy tweets database [28] contains tweets related to COVID vaccines from January 2021 to February 2023 and is available at doi.org/10.5281/zenodo.4526494. In this study we used tweets geolocated to US counties from February 6th to August 9th, 2021. The beginning of this window coincides with vaccines becoming widely available in the US, and the end of the window marks approximately the time when the vaccine uptake rate began to slow substantially. The primary features of this data are the text content, timestamps, and geolocations of the tweets. In addition, we used CDC records of COVID cases, deaths, and vaccinations in US counties [29], and Mønsted and Lehmann’s antivaccine tweets dataset [30], which supplements our own labeled data for training a tweet classifier. Additional details are provided in Supplementary Methods. Code is available at github.com/osome-iu/effects_of_antivax_tweets_on_covid19_outcomes.
@@ 5.2 Antivaccine Tweet Classifier
To track the volume of antivaccine content produced in each county, we built a text classifier that determines if a tweet expresses antivaccine sentiment. The classifier takes the text of a tweet as input and returns a label, either “antivaccine” or “other.” It was trained on 4,200 tweets labeled by two human annotators and 2,000 tweets labeled by Mønsted and Lehmann [30]. The classifier achieved good accuracy ($F_{1}=0.74$) on a hold-out set of 900 tweets, which were randomly sampled from our set of geolocated covid-related tweets. Further details about the data annotations, classification model, and results are found in Supplementary Methods.
@@ 5.3 Exposure to Antivaccine Tweets
To measure the impact of antivaccine tweets on people’s propensity to get vaccinated, we measure the amount of antivaccine Twitter content to which people are exposed at the county level. Intuitively, if people in county $i$ reshare a lot of content from people in county $j$, this implies that they are exposed to content from county $j$ — exposure is a necessary, though not sufficient, condition for sharing. Therefore we can use resharing of COVID-related content as proxy for exposure to such content. We further assume that exposure to anti-vaccine content is proportional to exposure to COVID-related content in general. Based on these assumptions, the per-capita antivaccine exposure rate in county $i$ at time $t$ is defined as
@@
where $N_{i}$ is the population of county $i$, $T_{j,t}$ is the number of antivaccine tweets in county $j$ during time window $t$, and $W_{ij}$ is the number of times that any COVID-related tweets posted by users in county $j$ were retweeted by users in county $i$ during our observation period. Here the retweet network is only used to estimate the strength of information flow between counties; our measure of exposure does not require that people retweet antivaccine content to be counted as exposed.
@@ 5.4 SIRVA Model
In the SIRVA epidemic model we developed for this work (see Figure 2), we make three simplifying assumptions: first, that vaccinated people ($V$) do not become infected; second, that vaccine-hesitant people ($A=\alpha S$) do not become vaccinated; and finally, we ignore the lag time between exposure to COVID and active infection.
The dynamic equations for SIRVA in each county can be written as:
@@
$\alpha_{t}=\alpha_{0}+\int_{t^{\prime}<t}\frac{d\alpha_{t^{\prime}}}{dt^{\prime}}\,dt^{\prime}=\alpha_{0}+\int_{t^{\prime}<t}\gamma(1-\alpha_{t^{\prime}})\,dt^{\prime}.$ (7)
Finally, we break the conversion rate $\gamma$ into components due to exposure to antivaccine tweets ($\gamma_{e}$) and other factors ($\gamma_{p}$): $\gamma=\gamma_{p}+\gamma_{e}E_{t}$, where $E_{t}$ is the antivaccine exposure rate in the county at time $t$. For our analyses we use a time interval of 8 days between data points.
We also define a simpler comparison model, SIRV, which is a special case of SIRVA where $\alpha_{t}=0$ for all $t$. This allows us to test how including vaccine hesitancy and its dynamics impacts the model’s predictive performance. Additionally, we ran a static version of SIRVA, SIRVA_static, in which the vaccine-hesitant population is static in each county, i.e., $\alpha_{t}=\alpha_{0}$ for all $t$, and thus $\gamma_{e}=\gamma_{p}=0$; we found this model has slightly worse model fit than SIRVA by the ELPD-LOO criterion.
Another possible model would include a “word of mouth” feedback term, $\gamma_{a}\alpha$, in $\gamma$, accounting for the endogenous spread of vaccine hesitancy within a county as a social contagion. We also tested a model with this dynamic effect and found similar results to the SIRVA model described here (i.e. $\gamma_{e}$ was of similar magnitude, as were its confidence intervals). We dropped this feature of the model for simplicity.
5.5 Effect Estimations
We use causal graphical modeling (see Supplementary Methods) to derive an expression for the average treatment effect (ATE) of exposure $E$ on the per capita vaccine uptake rate $\frac{1}{N_{i}}\dot{V}=\frac{1}{N_{i}}\frac{dV}{dt}$ from the dynamical equations of the SIRVA model (Eqs. 27):
$\text{ATE}=\frac{d}{dE_{t-1}}\mathbb{E}\Big{[}\frac{\dot{V_{t}}}{N_{i}}\mid\alpha_{t-1}\Big{]}\approx-\mathbb{E}\Big{[}\frac{1}{N_{i}}\nu_{t-1}\,S_{t-1,i}\;\gamma_{e}\,(1-\alpha_{t-1,i})\Big{]}$ (8)
where the expectation value is taken over times $t$, counties $i$, and the posterior distributions of $\nu$, $\gamma_{e}$, $\alpha$, and the derivation is detailed in Supplementary Methods (Eq. A3). This results in a figure of $3.2\times 10^{-4}$ (with a $95\%$ credible interval between $3.8\times 10^{-4}$ and $2.6\times 10^{-4}$) vaccinations per daily tweet of exposure.
Given the ATE from Eq. 8, we wish to calculate how many vaccinations were prevented in each county and nationwide. Assuming the total change in vaccine uptake rate is relatively small, we can estimate the number of vaccinations per capita prevented each day in each county as $\frac{1}{N_{i}}{\Delta V}_{t,i}\approx(\text{ATE})E_{t-1,i}$, and then use this estimation to calculate the total number of vaccines prevented nationwide as ${\Delta V}=N\,\Delta t\;\mathbb{E}_{i,t}[\frac{\Delta V_{t,i}}{N_{i}}]$, where $N$ is the total population of the US and $\Delta t$ is the length of the observation period (see Supplementary Methods).
We calculate the number of people across the whole US who remained unvaccinated as a result of exposure to antivaccine tweets to be 14,086 (see Supplementary Methods). We then wish to estimate how many cases and deaths may have been prevented among these people if they had been vaccinated instead. Assuming their infection and death rates were typical for unvaccinated people during the February-August 2021 time period (3,870 cases and 57 deaths per 100,000 people, see Supplementary Methods), we estimate that about 511 cases and 8 deaths occurred in this population. Of these cases and deaths, approximately 93% of cases and 94% of deaths may have been prevented by vaccination [31, 32], yielding the numbers of COVID cases and deaths attributable to antivaccine tweets, as reported in our results.
5.6 Model Selection Criteria
We can measure each model’s expected out-of-sample model performance using leave-one-out cross validation (LOO), approximated with Pareto-smoothed importance sampling (PSIS). The PSIS-LOO criterion [33] is designed to estimate out-of-sample predictive performance by approximating the expected log pointwise predictive density (ELPD) for a new dataset from the observed dataset without refitting the model to the data. The criterion is robust and efficient and represents the current state of the art for Bayesian model comparison. We therefore compare the performance of the SIRVA model to simpler models by measuring each model’s Bayesian LOO estimate of ELPD (ELPD-LOO). Values of ELPD-LOO closer to zero indicates better fit to the data. Metrics were computed on a sampled data set of 400 random counties and 24 evenly spaced dates spanning the observation period from February to August 2021.
@@ 5.6 Model Selection Criteria
Abbreviations
ATE: Average Treatment Effect. CDC: Centers for Disease Control and Prevention. CGM: Causal Graphical Model. ELPD: Expected Log Pointwise Predictive Density. HDI: High Probability Interval. LOO: Leave-One-Out. MCMC: Markov Chain Monte Carlo. PSIS: Pareto-Smoothed Importance Sampling. SIR: Susceptible-Infected-Recovered. SIRV: Susceptible-Infected-Recovered-Vaccinated. SIRVA: Susceptible-Infected-Recovered-Vaccinated-Antivaccine.
@@
Supplementary Methods are included in Appendix.
@@
Ethics approval
The collection of public posts from Twitter was deemed exempt from review by the Indiana University Institutional Review Board (protocol 1102004860).
Data and code availability
We use three main data sources in this work: the CoVaxxy Tweets database, which contains Tweets related to COVID vaccines [28] and is available at doi.org/10.5281/zenodo.4526494; CDC records of COVID cases, deaths, and vaccinations in US counties [29]; and Mønsted and Lehmann’s antivaccine tweets dataset [30], which supplements our own labeled data for training a tweet classifier. Details are provided in Supplementary Methods. Data and code are available at github.com/osome-iu/effects_of_antivax_tweets_on_covid19_outcomes.
Funding
This work was supported in part by the Knight Foundation and Craig Newmark Philanthropies.
Author contribution
All authors contributed to conceptualization of the work and data acquisition. JBr and Jbo developed the methodology. JBo led the investigation (software development, data curation and analysis, results validation, visualization) and wrote the original draft. JBr and FM provided supervision and reviewed and edited the manuscript. FM acquired lab funding and administered the project.
Acknowledgements
We are grateful to Marissa Donofrio for annotating tweets, to Bjarke Mønsted and Sune Lehmann for sharing additional annotation data, and to Vincent Jansen, Alessandro Flammini, and Yong-Yeol Ahn for helpful discussion. We also thank two anonymous reviewers who provided helpful feedback, highlighting an error in an earlier version of the manuscript.
@@ A.1 Data
The primary data source for this work is the CoVaxxy project [28] from Indiana University’s Observatory on Social Media. CoVaxxy collects tweets related to COVID vaccines and vaccine hesitancy, and geolocates the tweets to US counties when possible. The tweets were geolocated using Carmen [34], a tool which looks for geolocation meta data in tweets and self-identified locations in user profiles and post texts. Social bots were deliberately not excluded from our data, as bots may also contribute to population exposures to antivaccine content. Using this dataset, we can track the online discourse surrounding COVID vaccination in individual US counties.
The US Center for Disease Control (CDC) collected and published data on COVID health outcomes and vaccination in US counties, including cumulative cases, deaths, and vaccinations [29] for each county, for each day. This is the source of the public health metrics in this work. Due to differences in local reporting systems and reporting schedules, some entries in the cumulative counts are imputed by CDC as the last known value until they are updated with new reports from the counties or states. Counties within the state of Texas are excluded from our dataset because the official CDC data on COVID vaccinations does not contain information about Texas counties until after October 22, 2021.
To train our antivaccine tweet classifier, we require a dataset of tweets labeled as “antivaccine” or “other.” We supplement our own labeled data with a similar dataset created by Mønsted and Lehmann [30]. Restrictions apply to the availability of this dataset, which therefore is not publicly available. The data is, however, available from the authors upon reasonable request. Although this dataset is not specific to COVID antivaccine sentiment, we found its content to be similar enough to improve the performance of our classifier. We used the Mønsted and Lehmann dataset for training our model but not as part of test data used to evaluate our classifier’s performance.
@@ A.2 Antivaccine Tweet Classifier
To track the volume of antivaccine content produced in each county, we built a text classifier that determines if a tweet is expressing antivaccine sentiment. The classifier takes the text of a tweet as input and returns a label, either “antivaccine” or “other.” The classifier is a pre-trained neural network based on the RoBERTa language model [35], and is fine-tuned for our classification problem on a set of manually labeled tweets.
Our labeled training and test data were coded by two human annotators and examined for agreement. Cases where annotators disagreed were discarded. To be labeled as antivaccine, a tweet must “express the belief that safe, effective COVID vaccines are bad, ineffective, not actually a vaccine, or harmful (without specific evidence).” Most tweets expressing the belief that COVID vaccines are harmful made one of a few common claims, so these common claims were manually checked using reputable fact-checkers (e.g., PolitiFact, FactCheck.org) and CDC publications; in general, the common claims of harm were found to be false and labeled as antivaccine. We did not code other possible categories of vaccine sentiment, such as “pro-vaccination,” because our goal was to analyze the effect of antivaccine tweets.
The classifier was trained on 4,200 labeled tweets from the CoVaxxy dataset. To increase training data volume and variety, we also added to this training set 2,000 tweets randomly selected from the Mønsted and Lehmann dataset [30], which were labeled by three human annotators; cases where annotators disagreed were discarded. The model was tested on a hold-out set of 900 labeled tweets from the CoVaxxy dataset, labeled by the same method and annotators as the training data. We did not use cross-validation because (i) the training data includes Mønsted and Lehmann’s tweets, which do not match 2021 COVID-related tweets, and (ii) our sample of training tweets from the CoVaxxy dataset deliberately included a greater proportion of antivaccine tweets to help with classifier learning. Instead we produced the test dataset by pure random sampling from the CoVaxxxy data to ensure that classification metrics like $F_{1}$ are unbiased.
The classifier was evaluated using common classification metrics. It has an accuracy of $0.94\pm 0.02$, an $F_{1}$ of $0.74\pm 0.07$, and a Matthews Correlation Coefficient [36] of $0.72\pm 0.08$, where 95% confidence intervals are computed by the bootstrap procedure [37]. This classifier was used to determine the number of antivaccine tweets geolocated in each county on each date. While some individual tweets may be misclassified, we believe the performance of the classifier is adequate to determine population-level trends and relative magnitudes in the prevalence of antivaccine tweets.
@@ A.3 Estimating Parameters
Our goal is to infer the parameters of the SIRVA model from the data. In particular, we want to infer $\gamma_{e}$ to understand whether exposure to antivaccine tweets leads additional people to become unwilling to be vaccinated. To estimate the likely range of the model parameters, we use Bayesian inference with MCMC sampling [38]. The goal of Bayesian inference is to find the probability distribution (which we call the posterior distribution) that describes the likely values of a parameter given the data. From the posterior distributions, we can get a mean estimate, the 95% high probability interval (“HDI”), and a $p$-value for each parameter.
At a high level, Bayesian MCMC inference samples random parameter values proportionally to their likelihood given the data. The MCMC algorithm begins by sampling from a prior distribution, which is defined by specifying plausible ranges of the parameters, and gradually converges to the posterior distribution. We use the standard NUTS algorithm [39] implemented in the NumPyro package [40].
To specify a likelihood function for the SIRVA model given the data, we will define probability distributions for the daily changes in each of our observed variables: cumulative cases ($C=I+R$), vaccinations ($V$), susceptible individuals (S=N(V+C)S=N-(V+C)), and recovered individuals ($R$). We assume $R_{t}\approx C_{t-8}$, based on a typical time from initial symptoms to non-infectious status of 10 days [41] and a lag time between initial symptoms and a positive test of about 2 days.
@@
These distributions have high variance relative to the plausible parameter ranges so that the priors do not strongly influence the final inferred parameter values. We note that $\nu$ has a somewhat sharper prior than other parameters, informed by knowing the approximate number of total vaccinations in the studied time period. This sharper prior was necessary to eliminate identifiability issues between $\nu$ and $\gamma_{e}$.
For the vaccine-hesitancy parameters, $\gamma_{e}$ and $\gamma_{p}$, we choose priors centered at zero. This way, if the posterior is found to be non-zero, we know the prior did not bias that conclusion. For $\alpha_{0}$, we choose a weak prior based on general estimates of vaccine hesitancy in the population [42]. We use these priors:
@@
In our model, the parameters $\nu$ and $\beta$ can vary over time, following prior work [43, 44]. The $\nu$ parameter is allowed to change once every six weeks to account for changing nationwide vaccine availability. The $\beta$ parameter is allowed to change once every three weeks to account for changing mean reproduction numbers associated with the emergence of new variants (e.g., the Delta variant late in our observation window) and changing public health policies (e.g., the imposition or lifting of lockdowns and mask mandates). The multiple values of these parameters are inferred by the MCMC sampler just like the other parameters.
Our dataset comprises 1,319 counties and 188 dates for which we have sufficient Twitter and public health data. We use a subsample of this data to perform our inferences. We sample by date to minimize the effect of temporal autocorrelations in the data; specifically, we use every 8th day in the time series data for each county. This interval is chosen for two reasons: (i) it is slightly more than a week, so it smooths over data lags associated with weekly reporting cycles, and (ii) it is the approximate recovery time used to compute $R$. We also randomly sample by county to limit the number of model parameters (each county $i$ introduces an additional parameter $\alpha_{0,i}$). Our Bayesian MCMC inference tools struggle with numerical stability when the numbers of parameters and data points get too large. We therefore use 400 randomly selected counties and 24 evenly-spaced dates from our data. Although this sampling ultimately reduces the precision of our inferred parameter values, we do not believe it introduces any systematic biases; results are consistent across different random samples of the counties.